OpenStudy (anonymous):

How do I prove this identity: tanx + 1/tanx = 1/sinxcosx? o.O

OpenStudy (anonymous):

\[tanx + 1\div tanx = 1 \div sinxcosx\]

OpenStudy (saifoo.khan):

Try it. D;

OpenStudy (anonymous):

I got

OpenStudy (saifoo.khan):

@milliex51 , yay!

OpenStudy (anonymous):

\[sinx/cosx + 1/(sinx/cosx)\] \[1 x \cos ^{2}x/\sin ^{2}x = 1/sinxcosx\]

OpenStudy (anonymous):

and I'm stuck.. D:

OpenStudy (saifoo.khan):

which one is the main problem? Lol

OpenStudy (anonymous):

what do you mean?

OpenStudy (saifoo.khan):

You wrote two things above

OpenStudy (anonymous):

oh those are my steps..

OpenStudy (saifoo.khan):

WAIT. i did this same question some days ago! are you doing CIE by any chance?

OpenStudy (anonymous):

what's CIE? o.o

OpenStudy (saifoo.khan):

Cambridge Intl' Examinations.

OpenStudy (saifoo.khan):

That question was in my exam yesterday! OMG! im shocked.

OpenStudy (anonymous):

oh noooo

OpenStudy (anonymous):

LOL, sorry never did CIE

OpenStudy (saifoo.khan):

tanx + 1/tanx = 1/sinxcosx? \[\frac{\sin x}{\cos x} + \frac{cosx}{\sin x}\]Solve. :)

OpenStudy (anonymous):

is it uh... sin^2x / cos^2x...?

OpenStudy (saifoo.khan):

No, we will take "cosx sinx" as Lcm. \[\frac{\sin x(sinx) + cosx(\cos x)}{\cos x \sin x}\]

OpenStudy (anonymous):

so sin^2x + cos^2x / cosxsinx... but how? D:

OpenStudy (saifoo.khan):

Add, 1/6 + 1/7

OpenStudy (anonymous):

oh it's 7/42 + 6/42?

OpenStudy (saifoo.khan):

Yes but how you got it? you took LCM right? and then multiplied it by numerator..

OpenStudy (saifoo.khan):

We will do the same thing there.

OpenStudy (anonymous):

yes! thank you! :)

OpenStudy (saifoo.khan):

Welcome.